OpenStudy (anonymous):

Using calculus, it can be shown that if a ball is thrown upward with an initial velocity of 32 ft/s from the top of a building 594 ft high, then its height h above the ground t seconds later will be h = 594 + 32t − 16t2.During what time interval will the ball be at least 34 ft above the ground?

OpenStudy (anonymous):

Solve 594 + 32 t - 16 t^2=34 for t

OpenStudy (anonymous):

you get t=-5 and t=7 between 0 and 7

OpenStudy (anonymous):

\(a(t)=-32\) as gravity sucks \(v(t)=\int a(t)dt =-32t+v_0\) and since you are told \(v_0=32\) we get \(v(t)=-32t+32\) then \(h(t)=\int v(t)dt=-16t^2+32t+s(0)\) and since \(h(0)=594\) we have \[h(t)=-16t^2+32t+594\] i wish i could learn to read i thought it said "show" not "it can be shown" what a waste

OpenStudy (anonymous):

But not reading, you gave a review of how to get it.